# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 2.3 Structured multi-marginal optimal transport
— see this by noting that $\langle C,Q_{\lambda}\rangle+\varepsilon\langle Q_{\lambda},\log Q_{\lambda}\rangle-\sum_{j\in[m]}\lambda_{j}^{\top}P_{j}(Q_{\lambda})=0$. Inserting this solution $Q_{\lambda}$ into the Lagrangian function yields the dual functional $\varphi$. The dual of problem (9) becomes
$\max_{\lambda}\,\varphi(\lambda)=\max_{\lambda}\,\sum_{j\in[m]}\lambda_{j}^{\top}\upmu_{j}-\varepsilon\langle K,U_{\lambda}\rangle+\varepsilon n^{m+1}.$ (10)
@@
Note that the dual variables are unconstrained since the corresponding constraints of the primal problem are equality constraints (see, for example, [47]). One can show that for a primal-dual pair like the one above, strong duality holds (see, for example, [47]), meaning that the respective problem values of problems (9) and (10) coincide. Moreover, the optimal solution of the primal problem (9) is $Q_{\lambda^{*}}$, where $\lambda^{*}$ is the optimal solution of the dual problem (10). The optimal solution $Q_{\lambda^{*}}$ of the primal problem (9) can therefore be obtained by first solving the dual problem (10).
@@
Since the dual functional $\varphi$ is concave, the dual problem is a convex problem. Its maximisers are thus found as stationary points of $\varphi$. The optimal $\lambda$ therefore satisfies $0=\partial_{\lambda_{j}(i_{j})}\varphi(\lambda)=\upmu_{j}(i_{j})-P_{j}(K\odot U_{\lambda})(i_{j})$ for $i_{j}\in\{1,\dots,n\}$ and $j\in[m]$. In order to simplify the expressions, the solution is reparameterised in terms of $u=\{u_{j}\}_{j\in[m]}$ where $u_{j}:=\exp(\lambda_{j}/\varepsilon)$ for $j\in[m]$. It follows that the tensor $U_{\lambda}$ can be written $U_{\lambda}(i_{0},\dots,i_{m})=\prod_{\ell\in[m]}u_{\ell}(i_{\ell})$, for $i_{0},\dots,i_{m}\in\{1,\dots,n\}$. Since $U_{\lambda}$ is a rank one tensor, it holds that $P_{j}(K\odot U_{\lambda})=u_{j}\odot P_{j}(K\odot U_{\lambda}^{-j})$, where $U_{\lambda}^{-j}(i_{0},\dots,i_{m}):=\prod_{\ell\in[m]\backslash j}u_{\ell}(i_{\ell})$. The optimal $u$ therefore satisfies
$u_{j}=\upmu_{j}\oslash P_{j}(K\odot U_{\lambda}^{-j}),\quad j\in[m].$ (11)
@@
The computational bottleneck of the formula (11) is evaluating the projection, which requires manipulating an $n^{m+1}$-tensor. Consequently, the computational complexity increases exponentially as the number of marginals $m+1$ grows [44], and entropic regularisation alone is thus not enough for efficient computation of the problem. Fortunately, for many such multi-marginal problems the cost function has graph-structures that can be exploited in combination with entropic regularisation [2, 4, 8, 23, 30, 31, 33, 46]. These are referred to as graph-structured multi-marginal optimal transport problems. For cost functions that decouples sequentially, i.e. are of the form
$C(i_{0},\dots,i_{m})=\sum_{j\in[m]\backslash 0}C_{j}(i_{j-1},i_{j}),\quad i_{0},\dots,i_{m}=1,\dots,n$ (12)
@@
for some matrices $C_{1},\dots,C_{m}\in\mathbb{R}^{n\times n}$, the projection can be computed efficiently using matrix-vector multiplications. This is described in the following proposition. [^4] In [23], Proposition 2.8 was stated in the special case when the matrices $K_{j}$, $j\in[m]\backslash 0$, do not depend on the index $j$. However, it is straightforward to generalise it to the version given here.
@@
Let the elements of the tensors $K$ and $U$ be of the form
$K(i_{0},\dots,i_{m})=\prod_{j\in[m]\backslash 0}K_{j}(i_{j-1},i_{j})\qquad\mbox{ and }\qquad U(i_{0},\dots,i_{m})=\prod_{j\in[m]}u_{j}(i_{j})$
for matrices $K_{j}\in\mathbb{R}^{n\times n}$, $j\in[m]\backslash 0$, and vectors $u_{j}\in\mathbb{R}^{n}$, $j\in[m]$. Then,
$P_{j}\big{(}K\odot U\big{)}=$ $\big{(}u_{0}^{\top}K_{1}\operatorname{diag}(u_{1})K_{2}\dots K_{j-1}\operatorname{diag}(u_{j-1})K_{j}\big{)}^{\top}\odot u_{j}$
$\odot\big{(}K_{j+1}\operatorname{diag}(u_{j+1})K_{j+2}\dots K_{m-1}\operatorname{diag}(u_{m-1})K_{m}u_{m}\big{)},\quad j\in[m].$
@@ 3 Reformulation as a structured OT problem
We now show how each MOT problem of the form (6) can be identified with a structured multi-marginal OT problem whose optimal solution can be computed efficiently. This is carried out in two main steps, where the first one is to reduce the path dependency of the martingale constraint. Once this is done, we will transform the resulting problem to tensor form. A key insight throughout is to leverage on the sequential structure inherent in the problem.
@@ 3.1 Markovian reformulation of the problem
When comparing the form (5) of our class of payoff functions with the form (12) of the cost functions characterising the sequentially structured OT problems, we note that they are consistent. Unfortunately, the presence of the martingale constraint introduces a dependency of each marginal to all previous marginals, ruining the structure of the problem. Therefore, our class of MOT problems cannot directly be solved via the techniques mentioned in Section 2.3, despite the payoff function being of the appropriate form. It can be noted though that when the price process $S$ is a Markov process under the pricing measure $\mathbb{Q}$, the martingale constraint simplifies to only connecting adjacent marginals to each other. This recovers the structure of the problem, so that the projections can be computed efficiently. Leveraging on these insights, consider the modified multi-marginal OT problem
@@
where $X$ is given in equation 4. Note that since the constraints (13b) and (13c) are linear in $\mathbb{Q}$, the feasible set of problem (13) is convex. We now show that any optimal solution to problem (13) can be used to find an optimal solution to problem (6).
@@
The key idea behind the above result is that if a model $(\Omega,\mathcal{F},\mathbb{Q},S)$ satisfies equations (13b) and (13c), where the process $X$ is given by equation 4, then there exists a Markov process $(\tilde{S},\tilde{X})$, supported on some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, such that equations (13b) and (13c) hold, such that $(\tilde{S},\tilde{X})$ is of the form (4), and such that the corresponding value of the objective function (13a) is the same as for the original model $(\Omega,\mathcal{F},\mathbb{Q},S)$. In order to prove Theorem 3.1, we first establish a few intermediate results.
Fix any tuple $(\Omega,\mathcal{F},\mathbb{P},S)$, consisting of a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a stochastic process $S$ on that space, and define a process $X$ out of the process $S$ as in equation 4. Let $\nu:\mathcal{B}(\mathbb{R}^{2})\rightarrow[0,1]$ and $\kappa_{t}:\mathbb{R}^{2}\times\mathcal{B}(\mathbb{R}^{2})\rightarrow[0,1]$, for $t\in[T]\backslash 0$, be defined by
$\nu(A):=\mathbb{P}\left((S_{0},X_{0})\in A\right),\quad A\in\mathcal{B}(\mathbb{R}^{2})$
and
$\kappa_{t}(y,A):=\mathbb{P}\left((S_{t},X_{t})\in A|(S_{t-1},X_{t-1})=y\right),\quad A\in\mathcal{B}(\mathbb{R}^{2}),\quad y\in\mathbb{R}^{2},$
respectively, where we let the regular conditional distribution be equal to some real-valued constant if conditioning on a null event. Then $\nu$ is the initial distribution of the joint process $(S,X)$ on $\mathbb{R}^{2}$, while $\{\kappa_{t}\}_{t=1}^{T}$ is a family of kernels on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. Indeed, since for every $t\in[T]$ the random element $(S_{t},X_{t})$ takes values in $\mathbb{R}^{2}$ and $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$ is a Borel space, the regular conditional distribution given above exists and is almost everywhere equal to a kernel (see, for example, [41, Theorem 8.5]). The Kolmogorov extension theorem then guarantees the existence of an $\mathbb{R}^{2}$-valued Markov process $(\tilde{S},\tilde{X})$, defined on some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, with initial distribution $\nu$ and transition kernels $\{\kappa_{t}\}_{t=1}^{T}$ (see, for example, [41, Theorem 11.4]); that is,
~$\tilde{\mathbb{P}}((\tilde{S}_{0},\tilde{X}_{0})\in A)=\nu(A),\quad A\in\mathcal{B}(\mathbb{R}^{2})$ (14a)
~$\tilde{\mathbb{P}}((\tilde{S}_{t},\tilde{X}_{t})\in A|(\tilde{S}_{t-1},\tilde{X}_{t-1})=y)=\kappa_{t}(y,A),\quad A\in\mathcal{B}(\mathbb{R}^{2}),\quad y\in\mathbb{R}^{2},\quad t\in[T]\backslash 0.$ (14b)
It can be taken as the canonical process on the probability space given by $\tilde{\Omega}=(\mathbb{R}^{2})^{T+1}$, ~$\tilde{\mathcal{F}}=\sigma((\tilde{S}_{0},\tilde{X}_{0}),\dots,(\tilde{S}_{T},\tilde{X}_{T}))$ and ~$\tilde{\mathbb{P}}=\nu\otimes\kappa_{1}\otimes\dots\otimes\kappa_{T}$.
Lemma 3.2.
Let the stochastic processes $(S,X)$ and $(\tilde{S},\tilde{X})$ be as defined above. Then the law of $((\tilde{S}_{t-1},\tilde{X}_{t-1}),(\tilde{S}_{t},\tilde{X}_{t}))$ is equal to the law of $((S_{t-1},X_{t-1}),(S_{t},X_{t}))$ for all $t\in[T]\backslash 0$.
Proof.
It is immediate from equation 14a that $\text{Law}((\tilde{S}_{0},\tilde{X}_{0}))=\nu=\text{Law}((S_{0},X_{0}))$. Then fix any $t\in[T]\backslash 0$ and assume that
$\text{Law}((\tilde{S}_{t-1},\tilde{X}_{t-1}))=\text{Law}((S_{t-1},X_{t-1})).$
Let $F_{(\tilde{S}_{t-1},\tilde{X}_{t-1})}$ and $F_{(S_{t-1},X_{t-1})}$ denote the joint cumulative distribution functions corresponding to $\text{Law}((\tilde{S}_{t-1},\tilde{X}_{t-1}))$ and $\text{Law}((S_{t-1},X_{t-1}))$, respectively. Then $F_{(\tilde{S}_{t-1},\tilde{X}_{t-1})}=F_{(S_{t-1},X_{t-1})}$, so for any $A\in\mathcal{B}(\mathbb{R}^{2})$,
~$\tilde{\mathbb{P}}((\tilde{S}_{t},\tilde{X}_{t})\in A)$ $=\int_{\mathbb{R}^{2}}\kappa_{t}(y,A)\mathrm{d}F_{(\tilde{S}_{t-1},\tilde{X}_{t-1})}(y)$
$=\int_{\mathbb{R}^{2}}\kappa_{t}(y,A)\mathrm{d}F_{(S_{t-1},X_{t-1})}(y)=\mathbb{P}((S_{t},X_{t})\in A),$
where the second equality follows from equation 14b. Since the above holds for any $t\in[T]$, the assertion follows from this in combination with equation 14b. ∎
Remark 3.3.
In case the joint process $(S,X)$ is a Markov process under $\mathbb{P}$, it holds that
$\text{Law}((\tilde{S}_{0},\tilde{X}_{0}),\dots(\tilde{S}_{T},\tilde{X}_{T}))=\text{Law}((S_{0},X_{0}),\dots,(S_{T},X_{T})).$
We now show that the Markov process $(\tilde{S},\tilde{X})$ is almost surely of the form given in equation 4.
Lemma 3.4.
The process $(\tilde{S},\tilde{X})$ satisfies equation 4 almost surely, that is,
$\tilde{X}_{t}=\begin{cases}h_{0}(\tilde{S}_{0}),&t=0\\ h_{t}(\tilde{S}_{t-1},\tilde{S}_{t},\tilde{X}_{t-1}),&t=1,\dots,T\end{cases}\quad a.s.$
Proof.
The assertion for $t=0$ follows from Lemma 3.2 after noting that
$\big{\{}\tilde{X}_{0}=h_{0}(\tilde{S}_{0})\big{\}}=\big{\{}(h_{0}(\tilde{S}_{0}),\tilde{X}_{0})\in\{(x,x):x\in\mathbb{R}\}\big{\}}$
and similarly for the event $\{X_{0}=h_{0}(S_{0})\}$. Indeed,
~$\tilde{\mathbb{P}}((h_{0}(\tilde{S}_{0}),\tilde{X}_{0})\in\{(x,x):x\in\mathbb{R}\})$ =~$=\tilde{\mathbb{P}}((\tilde{S}_{0},\tilde{X}_{0})\in\{(h_{0}^{-1}(x),x):x\in\mathbb{R}\})$
$=\mathbb{P}((S_{0},X_{0})\in\{(h_{0}^{-1}(x),x):x\in\mathbb{R}\})$
$=\mathbb{P}((h_{0}(S_{0}),X_{0})\in\{(x,x):x\in\mathbb{R}\}),$
where the right-hand side is equal to one since by definition $X_{0}=h_{0}(S_{0})$. By similar arguments for $t\in[T]\backslash 0$, ~$\tilde{\mathbb{P}}(\tilde{X}_{t}=h_{t}(\tilde{S}_{t-1},\tilde{S}_{t},\tilde{X}_{t-1}))=\mathbb{P}(X_{t}=h_{t}(S_{t-1},S_{t},X_{t-1}))$. The assertion for $t\in[T]\backslash 0$ follows, which completes the proof. ∎
We are now ready to prove Theorem 3.1.
@@
$\mathbb{E}_{\mathbb{Q}}\big{[}S_{t}|\sigma(S_{t-1},X_{t-1})\big{]}$ $=\mathbb{E}_{\mathbb{Q}}\big{[}\mathbb{E}_{\mathbb{Q}}[S_{t}|\sigma(S_{0},\dots,S_{t-1})]|\sigma(S_{t-1},X_{t-1})\big{]}$
@@
Take any tuple $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ and define $X$ out of $S$ as given in equation 4. Then note that there exists an $\mathbb{R}^{2}$-valued Markov process $(\tilde{S},\tilde{X})$, supported on a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}})$, that has the same initial distribution and transition kernels as the joint process $(S,X)$ (cf. (14)). Since by Lemma 3.2 the law of $(\tilde{S}_{t-1},\tilde{S}_{t},\tilde{X}_{t-1},\tilde{X}_{t})$ under $\tilde{\mathbb{Q}}$ is equal to the law of $(S_{t-1},S_{t},X_{t-1},X_{t})$ under $\mathbb{Q}$ for every $t\in[T]\backslash 0$, it follows that the law of $\tilde{S}_{t}$ under $\tilde{\mathbb{Q}}$ is equal to the law of $S_{t}$ under $\mathbb{Q}$ for every $t\in[T]$. Therefore $\tilde{S}_{t}\sim_{\tilde{\mathbb{Q}}}\mu_{t}$ for $t\in\mathcal{T}$. By similar arguments,
$\mathbb{E}_{\tilde{\mathbb{Q}}}\big{[}\tilde{S}_{t}|\tilde{S}_{t-1}=s,\tilde{X}_{t-1}=x\big{]}=\mathbb{E}_{\mathbb{Q}}\big{[}S_{t}|S_{t-1}=s,X_{t-1}=x\big{]},\quad s,x\in\mathbb{R},\quad t\in[T]\backslash 0.$
Then note that by definition of $\mathfrak{F}^{\prime}$, the joint process $(S,X)$ satisfies $\mathbb{E}_{\mathbb{Q}}[S_{t}|S_{t-1}=s,X_{t-1}=x]=s$ for every $s,x\in\mathbb{R}$ and $t\in[T]\backslash 0$. Combining this with the above and with the Markov property of $(\tilde{S},\tilde{X})$ yields that $\tilde{S}$ is a martingale with respect to its own filtration under $\tilde{\mathbb{Q}}$. Therefore, $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})\in\mathfrak{F}$. Moreover, it follows from Lemma 3.2 that
@@
where the left-hand side is the payoff associated with the model $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})$ due to Lemma 3.4. We have thus shown that for any $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ we can find $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})\in\mathfrak{F}$ such that their corresponding values of the objective function are equal. This implies that the values of the two problems coincide.
@@
Therefore, in general $(\Omega,\mathcal{F},\mathbb{Q},S)\not\in\mathfrak{F}$ when $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$. Consequently $\mathfrak{F}\neq\mathfrak{F}^{\prime}$. From this it can be noted that $\mathfrak{F}$ can be characterised as the subset of tuples $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ such that equation 15 holds. We immediately realise that every tuple $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ that is such that $(S,X)$ is a Markov process under $\mathbb{Q}$ also belongs to $\mathfrak{F}$. On the other hand, note that every tuple in $\mathfrak{F}^{\prime}$ does not correspond to an $\mathbb{R}^{2}$-valued Markov process. On a related note, one can show that the multi-marginal MOT problem that is subject an additional Markov constraint is equivalent to problem (13) in the same way as we have shown that problem (6) is equivalent to problem (13). This means that there are in total three problems that are all equivalent to each other, given that the payoff function is of the form (5). This holds despite the fact that this third problem is not convex.
@@
If we know that the optimal solution to a MOT problem is such that the price process is a Markov process, the process $X$ can be omitted; see this by letting $X_{t}=S_{t}$ for $t\in[T]$ in the above results. This reduces the size of the problem. This holds in particular for all bi-marginal problems since a stochastic process defined over two discrete time points trivially is a Markov process. Along the same lines, we emphasise that MOT problems characterised by a payoff of the form $\phi(S_{0},\dots,S_{T})=\sum_{t\in[T]\backslash 0}\phi_{t}(S_{t-1},S_{t})$, $\phi_{t}:\mathbb{R}^{2}\rightarrow\mathbb{R}$, can be shown to be equivalent with an OT problem subject to the constraint $\mathbb{E}_{\mathbb{Q}}[S_{t}|\sigma(S_{t-1})]=S_{t-1}$ for $t\in[T]\backslash 0$.

### v2

@@ 2.3 Structured multi-marginal optimal transport
Inserting this solution $\mathbf{Q}^{\lambda}$ into the Lagrangian function yields the dual functional $\varphi$. The dual of problem (9) becomes
$\max_{\lambda}\,\varphi(\lambda)=\max_{\lambda}\,\sum_{t\in[T]}\lambda_{t}^{\top}m_{t}-\varepsilon\langle\mathbf{K},\mathbf{U}^{\lambda}\rangle$ (10)
@@
— see this by noting that $\langle\mathbf{C},\mathbf{Q}^{\lambda}\rangle+\varepsilon\langle\mathbf{Q}^{\lambda},\log\mathbf{Q}^{\lambda}\rangle-\sum_{t\in[T]}\lambda_{t}^{\top}P_{t}(\mathbf{Q}^{\lambda})=0$. Note that the dual variables are unconstrained since the corresponding constraints of the primal problem are equality constraints (see, for example, [53]). One can show that for a primal-dual pair like the one above, strong duality holds (see, for example, [53]), meaning that the respective problem values of problems (9) and (10) coincide. Moreover, the optimal solution of the primal problem (9) is $\mathbf{Q}^{\lambda^{*}}$, where $\lambda^{*}$ is the optimal solution of the dual problem (10). The optimal solution $\mathbf{Q}^{\lambda^{*}}$ of the primal problem (9) can therefore be obtained by first solving the dual problem (10).
@@
Since the dual functional $\varphi$ is concave, the dual problem is a convex problem. Its maximisers are thus found as stationary points of $\varphi$. The optimal $\lambda$ therefore satisfies $0=\partial_{\lambda_{t}(i_{t})}\varphi(\lambda)=m_{t}(i_{t})-P_{t}(\mathbf{K}\odot\mathbf{U}^{\lambda})(i_{t})$ for $i_{t}\in\{1,\dots,n\}$ and $t\in[T]$. In order to simplify the expressions, the solution is reparameterised in terms of $u=\{u_{t}^{\lambda_{t}}\}_{t\in[T]}$ where $u_{t}^{\lambda_{t}}:=\exp(\lambda_{t}/\varepsilon)$ for $t\in[T]$. It follows that the tensor $\mathbf{U}^{\lambda}$ can be written $\mathbf{U}^{\lambda}(i_{0},\dots,i_{T})=\prod_{\ell\in[T]}u^{\lambda_{\ell}}_{\ell}(i_{\ell})$, for $i_{0},\dots,i_{T}\in\{1,\dots,n\}$. Since $\mathbf{U}^{\lambda}$ is a rank one tensor, it holds that $P_{t}(\mathbf{K}\odot\mathbf{U}^{\lambda})=u_{t}^{\lambda_{t}}\odot P_{t}(\mathbf{K}\odot\mathbf{U}^{\lambda}_{-t})$, where $\mathbf{U}^{\lambda}_{-t}(i_{0},\dots,i_{T}):=\prod_{\ell\in[T]\backslash t}u^{\lambda_{\ell}}_{\ell}(i_{\ell})$. The optimal $u^{\lambda}$ therefore satisfies
$u_{t}^{\lambda_{t}}=m_{t}\oslash P_{t}(\mathbf{K}\odot\mathbf{U}^{\lambda}_{-t}),\quad t\in[T].$ (11)
@@
The computational bottleneck of the formula (11) is evaluating the projection, which requires manipulating an $n^{T+1}$-tensor. Consequently, the computational complexity increases exponentially as the number of marginals $T+1$ grows [50], and entropic regularisation alone is thus not enough for efficient computation of the problem. Fortunately, for many such multi-marginal problems the cost function has graph-structures that can be exploited in combination with entropic regularisation [2, 4, 9, 26, 34, 35, 37, 52]. These are referred to as graph-structured multi-marginal optimal transport problems. For cost functions that decouples sequentially, i.e. are of the form
$\mathbf{C}(i_{0},\dots,i_{T})=\sum_{t\in[T]\backslash 0}C_{t}(i_{t-1},i_{t}),\quad i_{0},\dots,i_{T}=1,\dots,n$ (12)
@@
for some matrices $C_{1},\dots,C_{T}\in\mathbb{R}^{n\times n}$, the projection can be computed efficiently using matrix-vector multiplications. This is described in the following proposition. [^4] In [26], Proposition 2.9 was stated in the special case when the matrices $K_{t}$, $t\in[T]\backslash 0$, do not depend on the index $t$. However, it is straightforward to generalise it to the version given here.
@@
Let the elements of the tensors $\mathbf{K}$ and $\mathbf{U}$ be of the form
$\mathbf{K}(i_{0},\dots,i_{T})=\prod_{t\in[T]\backslash 0}K_{t}(i_{t-1},i_{t})\qquad\mbox{ and }\qquad\mathbf{U}(i_{0},\dots,i_{T})=\prod_{t\in[T]}u_{t}(i_{t})$
for matrices $K_{t}\in\mathbb{R}^{n\times n}$, $t\in[T]\backslash 0$, and vectors $u_{t}\in\mathbb{R}^{n}$, $t\in[T]$. Then,
$P_{t}\big{(}\mathbf{K}\odot\mathbf{U}\big{)}=$ $\big{(}u_{0}^{\top}K_{1}\operatorname{diag}(u_{1})K_{2}\dots K_{t-1}\operatorname{diag}(u_{t-1})K_{t}\big{)}^{\top}\odot u_{t}$
$\odot\big{(}K_{t+1}\operatorname{diag}(u_{t+1})K_{t+2}\dots K_{T-1}\operatorname{diag}(u_{T-1})K_{T}u_{T}\big{)},\quad t\in[T].$
@@ 3 Reformulation as a structured OT problem
We now show how an MOT problem of the form (6) can be identified with a structured multi-marginal OT problem whose optimal solution can be computed efficiently. This is carried out in two steps, where the first one is to reduce the path dependency of the martingale constraint. Once this is done, we will transform the resulting problem to tensor form. A key insight throughout is to leverage on the sequential structure inherent in the problem.
@@ 3.1 Markovian reformulation of the problem
When comparing the form (5) of our class of payoff functions with the form (12) of the cost functions characterising the sequentially structured OT problems, we note that they are consistent. However, the presence of the martingale constraint introduces a dependency of each marginal to all previous marginals, ruining the structure of the problem; consequently, our class of MOT problems cannot directly be solved via the techniques described in Section 2.3, despite the payoff function being of the appropriate form. Instead we consider the modified multi-marginal OT problem
@@
where $X$ is given in equation 4. Note that since the objective function (13a) and the constraints (13b) and (13c) are linear in $\mathbb{Q}$, the above is a convex problem. The key idea is that for every model $(\Omega,\mathcal{F},\mathbb{Q},S)$ that satisfies equations (13b) and (13c), where the process $X$ is given by equation 4, there exists a Markov process $(\tilde{S},\tilde{X})$, supported on some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}})$, with similar properties as the joint process $(S,X)$; in particular, this $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})$ is feasible to both problems (6) and (13) and it yields the same value of the objective function as the model $(\Omega,\mathcal{F},\mathbb{Q},S)$.
We now show that any optimal solution to problem (13) can be used to find an optimal solution to problem (6).
@@
Take any tuple $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ and define $X$ out of $S$ as given in equation 4. Then note that there exists an $\mathbb{R}^{2}$-valued Markov process $(\tilde{S},\tilde{X})$, supported on some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}})$, that has the same initial distribution and conditional distributions as the joint process $(S,X)$. Indeed, let $\nu:\mathcal{B}(\mathbb{R}^{2})\rightarrow[0,1]$ and $\kappa_{t}:\mathbb{R}^{2}\times\mathcal{B}(\mathbb{R}^{2})\rightarrow[0,1]$, for $t\in[T]\backslash 0$, be defined by $\nu(A):=\mathbb{Q}((S_{0},X_{0})\in A)$ and $\kappa_{t}(y,A):=\mathbb{Q}((S_{t},X_{t})\in A|(S_{t-1},X_{t-1})=y)$ for $A\in\mathcal{B}(\mathbb{R}^{2})$ and $y\in\mathbb{R}^{2}$. Since $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$ is a Borel space, the regular conditional distribution given above exists and is almost everywhere equal to a kernel (see, for example, [47, Theorem 8.5]). The Kolmogorov extension theorem then guarantees the existence of an $\mathbb{R}^{2}$-valued Markov process $(\tilde{S},\tilde{X})$, defined on some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}})$ [^5] It can be taken as the canonical process on the probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}})$ given by $(\tilde{\Omega},\tilde{\mathcal{F}})=(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))^{T+1}$ and ~$\tilde{\mathbb{Q}}=\nu\otimes\kappa_{1}\otimes\dots\otimes\kappa_{T}$., with initial distribution $\nu$ and transition kernels $\{\kappa_{t}\}_{t\in[T]\backslash 0}$ (see, for example, [47, Theorem 11.4]). It follows that
$\text{Law}((\tilde{S}_{t-1},\tilde{X}_{t-1}),(\tilde{S}_{t},\tilde{X}_{t}))=\text{Law}((S_{t-1},X_{t-1}),(S_{t},X_{t})),\quad t\in[T]\backslash 0.$ (14)
Indeed, it is immediate that $\text{Law}((\tilde{S}_{0},\tilde{X}_{0}))=\nu=\text{Law}((S_{0},X_{0}))$. Combining this with the definition of the kernel $\kappa_{1}$ yields that $\text{Law}((\tilde{S}_{0},\tilde{X}_{0}),(\tilde{S}_{1},\tilde{X}_{1}))=\text{Law}((S_{0},X_{0}),(S_{1},X_{1}))$, from where it follows that $\text{Law}((\tilde{S}_{1},\tilde{X}_{1}))=\text{Law}((S_{1},X_{1}))$. Equation 14 then follows by induction. Consequently, the law of $\tilde{S}_{t}$ under $\tilde{\mathbb{Q}}$ is equal to the law of $S_{t}$ under $\mathbb{Q}$ for every $t\in[T]$ and therefore $\tilde{S}_{t}\sim_{\tilde{\mathbb{Q}}}\mu_{t}$ for $t\in\mathcal{T}$. By similar arguments,
$\mathbb{E}_{\tilde{\mathbb{Q}}}\big{[}\tilde{S}_{t}|\tilde{S}_{t-1}=s,\tilde{X}_{t-1}=x\big{]}=\mathbb{E}_{\mathbb{Q}}\big{[}S_{t}|S_{t-1}=s,X_{t-1}=x\big{]}=s,\quad s,x\in\mathbb{R},\quad t\in[T]\backslash 0.$
Combining this with the Markov property of $(\tilde{S},\tilde{X})$ yields that $\tilde{S}$ is a martingale with respect to its own filtration under $\tilde{\mathbb{Q}}$. Therefore, $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})\in\mathfrak{F}$. Moreover, it follows from equation 14 that
@@
where the left-hand side is the payoff associated with the model $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})$ since the process $(\tilde{S},\tilde{X})$ satisfies equation 4 almost surely. Indeed, recall that the functions $h_{t}$ for $t\in[T]$ are measurable and hence it follows from equation 14 that ~$\tilde{\mathbb{Q}}(\tilde{X}_{t}=h_{t}(\tilde{S}_{t},\tilde{S}_{t-1},\tilde{X}_{t-1}))=\mathbb{Q}(X_{t}=h_{t}(S_{t},S_{t-1},X_{t-1}))$ for $t\in[T]\backslash 0$, where the right-hand side equals 1, and similarly for the case when $t=0$. We have thus shown that for any $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ we can find $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{Q}},\tilde{S})\in\mathfrak{F}$ such that their corresponding values of the objective function are equal. This implies that the values of the two problems coincide.
@@
Therefore, in general $(\Omega,\mathcal{F},\mathbb{Q},S)\not\in\mathfrak{F}$ when $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$. Consequently $\mathfrak{F}\neq\mathfrak{F}^{\prime}$. From this it can be noted that $\mathfrak{F}$ can be characterised as the subset of tuples $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ such that equation 15 holds. We immediately realise that every tuple $(\Omega,\mathcal{F},\mathbb{Q},S)\in\mathfrak{F}^{\prime}$ that is such that $(S,X)$ is a Markov process under $\mathbb{Q}$ also belongs to $\mathfrak{F}$. On the other hand, note that every tuple in $\mathfrak{F}^{\prime}$ does not correspond to an $\mathbb{R}^{2}$-valued Markov process. However, one can show that the multi-marginal MOT problem that is subject to such an additional Markov constraint is equivalent to problem (13) in the same way as we have shown that problem (6) is equivalent to problem (13); this means that there are in total three problems that are all equivalent to each other, given that the payoff function is of the form (5). We choose to solve problem (13) which is convex, in contrast to the problem with an explicit Markov constraint.
@@
If we know that the optimal solution to an MOT problem is such that the price process is a Markov process, then the process $X$ can be omitted; see this by letting $X_{t}=S_{t}$ for $t\in[T]$ in the above results. This reduces the size of the problem. This holds in particular for all bi-marginal problems since a stochastic process defined over two discrete time points trivially is a Markov process. Along the same lines, we emphasise that MOT problems characterised by a payoff of the form $\phi(S_{0},\dots,S_{T})=\sum_{t\in[T]\backslash 0}\phi_{t}(S_{t-1},S_{t})$, $\phi_{t}:\mathbb{R}^{2}\rightarrow\mathbb{R}$, can be shown to be equivalent to an OT problem subject to the constraint $\mathbb{E}_{\mathbb{Q}}[S_{t}|\sigma(S_{t-1})]=S_{t-1}$ for $t\in[T]\backslash 0$.